OpenStudy (anonymous):

What is the excluded value of x for x^2+3x+2 ---------- x^2+2x+1

OpenStudy (anonymous):

Excluded values make the denominator 0. You want to set the denominator equal to 0 and then solve that equation by factoring

OpenStudy (anonymous):

So would I subtract 2x from the top then?

OpenStudy (anonymous):

No, you don't have to do anything to the top. The equation to solve is x² + 2x + 1 = 0

OpenStudy (anonymous):

I have a feeling that I'm doing this wrong then. I got x^2+2x+1=0; 2x+2x+1=0; 4x+1=0; and then I subtract one from both sides; and I get 4x=-1; x=-4

OpenStudy (anonymous):

Why did you turn x² into 2x?

OpenStudy (anonymous):

Don't you multiply the square root by the number it is rooting from?

OpenStudy (anonymous):

no. and we can't take a square root of just one term. You'd have to do it for the whole equation.

OpenStudy (anonymous):

Do you want to try the quadratic equation to solve it?

OpenStudy (anonymous):

I just tried the equation and got -4 for the answer?

OpenStudy (anonymous):

have you ever done FOIL or reverse FOIL? Does this look familiar?